### v1

# VIRTUE: Visual-Interactive Text-Image Universal Embedder

Abstract
Multimodal representation learning models have demonstrated successful operation across complex tasks, and the integration of vision-language models (VLMs) has further enabled embedding models with instruction-following capabilities. However, existing embedding models lack visual-interactive capabilities to specify regions of interest from users (e.g., point, bounding box, mask), which have been explored in generative models to broaden their human-interactive applicability. Equipping embedding models with visual interactions not only would unlock new applications with localized grounding of user intent, which remains unexplored, but also enable the models to learn entity-level information within images to complement their global representations for conventional embedding tasks. In this paper, we propose a novel Visual-InteRactive Text-Image Universal Embedder (VIRTUE) that extends the capabilities of the segmentation model and the vision-language model to the realm of representation learning. In VIRTUE, the segmentation model can process visual prompts that pinpoint specific regions within an image, thereby enabling the embedder to handle complex and ambiguous scenarios more precisely. To evaluate the visual-interaction ability of VIRTUE, we introduce a large-scale Segmentation-and-Scene Caption Retrieval (SCaR) benchmark comprising 1M samples that aims to retrieve the text caption by jointly considering the entity with a specific object and image scene. VIRTUE consistently achieves a state-of-the-art performance with significant improvements across 36 universal MMEB (3.1%–8.5%) and five visual-interactive SCaR (15.2%–20.3%) tasks111The code, model, and benchmark will be released upon acceptance..
1 Introduction
Embedding models have recently transitioned from two-tower architectures (e.g., CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), SigLIP (Zhai et al., 2023)), which have been used for embedding-based evaluation (Girdhar et al., 2023) and cross-modal similarity matching (Hao et al., 2023; Han et al., 2024), to vision-language model (VLM)-based frameworks (e.g., GME (Zhang et al., 2025), LamRA (Liu et al., 2025b)) owing to VLMs’ ability to ingest arbitrary combinations of textual and visual inputs into a single embedding space. Thanks to their inherent instruction-following capabilities, adopting VLMs as embedding models generalizes effectively across a wide range of zero-shot multimodal reasoning applications, including interactive information retrieval (Jiang et al., 2025b) and retrieval-augmented generation (Liu et al., 2025a).
Although VLM-based embedding models support interactive use, they only accept text as the human-machine interaction modality. In contrast, visual prompts, which have recently attracted attention in generative applications (You et al., 2024; Yuan et al., 2024; Lian et al., 2025), serve as an important but overlooked interaction channel. Visual prompting can not only enhance the downstream generation performance (Li et al., 2022b) but also provide precise spatial localization for fine-grained understanding (Liu et al., 2024c). This is particularly advantageous for embedding-based tasks, as it allows models to respond to visual inputs from the user beyond traditional global matching by capturing entity-level cues, thereby improving retrieval precision and alignment while complementing global representations.
Considering visual-interactive image-to-image (I2I) and image-to-text (I2T) retrieval scenarios, as shown in Fig. 1, where a user aims to retrieve different entities within the same image but under a shared global context, current embedding models rely solely on holistic image representations and fail to leverage explicit visual interactive signals (e.g., bounding boxes, points, and masks provided by users). As a result, they cannot isolate and retrieve the targeted entity while maintaining awareness of the broader scene (e.g., “grass” for the I2I scenario and “on a table with desserts” for the I2T retrieval). One possible strategy is to convert visual prompts into textual descriptions to guide retrieval; however, embedding models are not trained with spatially grounded supervision, which limits their ability to generalize to such interactive tasks. Another intuitive approach is to crop the region of interest (Subramanian et al., 2022), which can improve fine-grained understanding but sacrifices global contextual cues for compositional reasoning (e.g., understanding an object within the full scene as presented in Appendix E.4.1.). This limitation gives rise to a central challenge: How can visual interaction capabilities be incorporated into embedding models, and how can we systematically evaluate their compositional reasoning on targeted image regions?
In this paper, we propose VIRTUE, a visual-interactive text-image universal embedder that combines an off-the-shelf segmentation model (SAM2 (Ravi et al., 2025)) with a pretrained VLM to jointly encode entity- and global-level information from images and the textual descriptions. For visual-interactive scenarios, VIRTUE processes user-provided visual prompts by the prompt encoder within the segmentation model; for non-interactive scenarios, the prompt encoder is fed uniformly sampled points to produce a feature map composed of multiple entity-level information. The VLM then ingests arbitrary combinations of image and text embeddings, where each image embedding comprises both an entity-level embedding (from the segmentation model) and a global image embedding (from the VLM’s vision encoder), and produces a single unified embedding for contrastive learning. In this manner, VIRTUE enables training on visual-interactive and non-visual-interactive data and supports entity-aware retrieval while preserving global scene context.
Since no existing benchmark evaluates visual-interactive embedding capabilities, we introduce
SCaR, a large-scale Segmentation-and-Scene Caption Retrieval benchmark for visual-interactive image-to-text retrieval.
In SCaR, an image together with a region of interest serves as a query, and the task is to retrieve the caption that describes the specified object in its global scene context.
We constructed SCaR from five publicly available datasets: RefCOCO+ (Yu et al., 2016), RefCOCOg Mao et al. (2016), VisualGenome (Krishna et al., 2017), COCO-Stuff (Caesar et al., 2018), and ADE20k (Zhou et al., 2017).
The annotations include images, bounding boxes, and captions that describe entities, relations, and the global scene context.
To increase difficulties in reasoning, negative distractors are generated by replacing one of three elements of the ground-truth caption via prompting GPT-4V (OpenAI, 2023) instead of random sampling; for datasets that lack human captions (e.g., ADE20k), we generated ground-truth captions via carefully designed prompts to GPT-4V.
To this end, SCaR comprises a vast collection of 1M samples that are divided into training and validation sets.
A distinguishing characteristic of the proposed SCaR dataset is its ability to evaluate not only visual-interactive reasoning but also compositional scenarios, requiring models to perform fine-grained, context-aware cross-modal reasoning that goes beyond global image matching.
In summary, our contributions are three-fold:
Method Novelty: We propose VIRTUE, a visual-interactive text-image universal embedder consisting of a VLM as well as a segmentation model to enable the visual interaction modality for human interactions. The segmentation model allows users to optionally provide different types of visual prompts via its prompt encoder and reinforces VIRTUE to capture entity-level representations in addition to global context.
Benchmark Novelty: As there is no public visual-interactive embedding benchmark, we introduce SCaR, composed of 1M samples for visual-interactive image-to-text retrieval, to evaluate VIRTUE’s capabilities. SCaR enables evaluation of advanced reasoning and compositional tasks in multimodal, visual-interaction-aware embedding scenarios that remain unexplored.
Experiment Novelty: VIRTUE outperforms state-of-the-art embedding models on 36 MMEB tasks with significant gains from 3.1% to 8.5% and achieves improvements of 15.2% to 20.3% on five SCaR tasks, showing that equipping embedding models with visual-interactive capabilities benefits both visual-interactive and non-visual-interactive scenarios.
2 Related Works
Multimodal Representation Learning. Early progress in text-image representation learning was driven by two-tower contrastive models that learn a joint embedding space by aligning an image encoder with a text encoder (e.g., CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), SigLIP (Zhai et al., 2023), OpenCLIP (Cherti et al., 2023)). These models provide effective global image-text matching and have served as foundation models for building vision-language models (VLMs) that tackle zero-shot downstream tasks (Tong et al., 2024; Wang et al., 2025a; Deitke et al., 2025). Subsequent work has advanced embedding performance along various dimensions. For instance, UniIR (Wei et al., 2024) finetunes CLIP/BLIP with the late fusion of text and image embeddings, and UniME (Gu et al., 2025) distills text knowledge from an LLM followed by two-stage negatives contrastive learning. Magiclens (Zhang et al., 2024) incorporates open-ended instructions into dual-encoder architectures with training on large-scale instruction datasets. More recently, VLMs that accept arbitrary mixtures of visual and textual inputs and are trained with instruction-style objectives have emerged as flexible and unified embedding providers that better perform on multimodal reasoning and compositional queries, e.g., E5-V (Jiang et al., 2024), VLM2Vec (Jiang et al., 2025b), GME (Zhang et al., 2025), and LamRA (Liu et al., 2025b). Despite these advancements, existing embedding models only support textual instructions and lack native support for direct visual prompts; they are typically trained only on holistic image-text alignment (Oh et al., 2024). In contrast, VIRTUE integrates a segmentation model with a pretrained VLM to fuse segmentation-derived, prompt-conditioned entity representations with global representations, producing unified embeddings that are both entity-aware and context-preserving, as evaluated on both MMEB and our newly constructed SCaR.
Interactive Embedding Benchmarks. Since embedding models have shifted from uni-modal matching benchmarks (e.g., BEIR (Thakur et al., 2021), MTEB (Muennighoff et al., 2023)) to instruction-based cross-modal global matching, recent studies have introduced instruction-based multimodal benchmarks to probe the reasoning abilities of embedding models. M-BEIR (Wei et al., 2024) is a multimodal retrieval benchmark encompassing eight tasks from diverse domains using text instructions, while MMEB (Jiang et al., 2025b) extends the evaluation to 36 multimodal datasets covering classification, VQA, retrieval, and grounding tasks to assess instruction-following across multifaceted perspectives. While we utilize MMEB to evaluate our proposed method for universal embedding abilities, these benchmarks focus on text-based instructions and do not evaluate visual-interactive scenarios in which visual prompts are provided as inputs. To fill this gap, we introduce SCaR, a large-scale interactive image-to-text retrieval benchmark where each query consists of an image as well as a target bounding box, and the task is to retrieve captions that describe the specified entity within its global scene context. SCaR is composed of five visual-grounding and referring-expression datasets that test caption retrieval for a region-in-context, with negative distractors generated by GPT-4V to stress-test entity-in-context discrimination beyond the simple random negatives used in MMEB.
3 SCaR: Segmentation-and-Scene Caption Retrieval Benchmark
3.1 SCaR Overview
Current publicly available benchmarks primarily evaluate text instruction-following capabilities for embedding models.
Although MMEB contains out-of-domain visual grounding tasks for RefCOCO (Kazemzadeh et al., 2014), it simplifies them by cropping the specified region as the target, thereby neglecting the broader scene context within an image.
Consequently, existing embedding models struggle with inputs that include visual regions of interest in visual-interactive retrieval tasks.
To address this limitation, we introduce
SCaR, a segmentation-and-scene caption retrieval benchmark that challenges models with reasoning and compositionality for text caption retrieval based on a given image and a specified region.
SCaR comprises images, segmentations with bounding boxes, and text captions from five public datasets: RefCOCO+ (Yu et al., 2016), RefCOCOg (Mao et al., 2016), VisualGenome (Krishna et al., 2017), COCO-Stuff (Caesar et al., 2018), and ADE20k (Zhou et al., 2017).
Distinct from existing benchmarks, SCaR provides multiple negative candidates per sample, which are generated by GPT-4V (OpenAI, 2023) through element-swapping in the ground-truth caption with false replacements, forming a large-scale benchmark of 1M samples covering diverse applicability.
Task Definition. The main difference between conventional image-to-text retrieval (e.g., MSCOCO_i2t in MMEB) and visual-interactive image-to-text retrieval is the additional region-of-interest input. Formally, given an input image and a bounding box with ten candidates , the goal is to find the most relevant text caption , where is the embedding model and sim denotes cosine similarity. Since SCaR requires models to reason about both the specified object and the broader scene, candidate captions are intentionally challenging and demand joint reasoning over fine-grained bounding-box details and global image context.
SCaR. We adopt GPT-4V to generate missing elements for the ground-truth caption as well as negative candidates.
Collected samples (left) are filtered via LLM-then-human inspection (right) to ensure quality. Each SCaR sample contains an image with a bounding box, one ground-truth caption, and nine distractors.
3.2 Collection Pipeline
Fig. 2 illustrates the collection pipeline to construct SCaR, where all datasets are first converted to the COCO format to enable unified processing. To balance evaluation efficiency and coverage, we randomly sample up to five objects per image. For each sample, the image size, original caption, object category, and bounding box222While some datasets provide segmentation masks, we observed that GPT-4V struggles to interpret them reliably. In contrast, bounding boxes are easier for GPT-4V to parse and align with prior prompt-based collection strategies (Wang et al., 2025b; Jiang et al., 2025a); therefore, we opt for bounding boxes for SCaR. are provided in the prompt template, as shown in Fig. 4. The prompt instructs GPT-4V to return a JSON object containing the ground-truth caption and nine negatives. Since some datasets lack complete descriptions in terms of <object> <relation> <scene> (e.g., ADE20k only provides object names, and some RefCOCO+ samples do not contain global scene context), the prompt asks GPT-4V to verify whether the caption contains all three elements, and to supplement any missing elements with careful and image-grounded descriptions.
Negative candidates are then generated via element-swapping from the gold caption. We define three swap strategies, each producing three candidates. 1) Global scene swap: Replace the scene phrase with a clearly distinct environment (e.g., from “table” to “picnic blanket”). 2) Relation swap: Modify the relation of the specified object by borrowing from nearby objects or introducing hallucinated interactions (e.g., from “on the table” to “tucked beside the glass” or “under the napkin”). 3) Object swap: Substitute the target object with another object, either plausible but incorrect within the image or fabricated (e.g., from “salad fork” to “wine glass” or “butter knife”). In early trials, we observed that GPT-4V often generated ambiguous variants (e.g., “traffic light” vs. “stoplight”, “zoo” vs. “safari park”), which could reduce the quality and reliability of the benchmark. To mitigate this, we explicitly add constraints to the instruction requiring swapped objects and scenes to belong to clearly distinct categories. We also encourage GPT-4V to produce diverse and creative negatives, ensuring coverage across different objects, relations, and scenes, while reducing the risk of overfitting to narrow patterns.
| Train | Evaluation | |||
| #Images | #Annotations | #Images | #Annotations | |
| RefCOCOg | 21,730 | 40,674 | 1,300 | 1,539 |
| RefCOCO+ | 16,847 | 38,807 | 1,500 | 2,764 |
| COCO-Stuff | 118,287 | 426,379 | 4,999 | 17,903 |
| VisualGenome | 86,414 | 357,583 | 5,000 | 15,571 |
| ADE20K | 20,210 | 94,271 | 2,000 | 9,368 |
| Total | 309,278 | 957,714 | 14,799 | 47,145 |
LLM-then-Human Inspection. Despite carefully designed prompts for guiding GPT-4V, some deviations from established rules result in the generation of subpar ground-truth and candidate generations, including generated ground-truth captions that lack scene context and negatives that involve synonymy or ambiguous objects and scenes. Therefore, we design a multi-stage filtering pipeline combining heuristic rules, LLM-based verification, and human inspection to improve dataset reliability and remove unethical samples. After the collection, GPT-4V is guided to verify whether each ground-truth caption contains all three elements <object> <relation> <scene> and to output these elements in a structured JSON format for both ground-truth and negative captions, as shown in Fig. 5. Then, WordNet (Miller, 1992) is applied to detect if any negative elements are synonyms of the corresponding gold elements. The sample is immediately discarded if any verification step fails. For the evaluation set, two independent human inspectors review all remaining samples with a focus on ambiguity, and remove any that are flagged by either inspector. For the training set, we assess quality indirectly by training models on it and evaluating them on the SCaR evaluation set (Sec. 5.3). This meticulous filtering regimen ensures the integrity and trustworthiness of the SCaR dataset, which comprises 957k training and 47k evaluation samples (see Tab. 1). More statistics and examples are provided in Appendix D.
4 VIRTUE
As illustrated in Fig. 3, since our goal is to create a general-purpose vision-language embedding model with the visual interaction capability, we propose VIRTUE, which consists of a VLM that processes visual and textual inputs as vision and language embeddings, respectively, and a segmentation model along with a segmentation-language connector that converts visual prompts (i.e., bounding boxes, masks, points, or sampled points when not explicitly provided) into the segmentation embeddings. Afterwards, the large language model (LLM) consumes the sequence of segmentation, vision, and language embeddings to generate the final input embedding using the final hidden state of the last token for contrastive learning.
4.1 Converting Visual and Textual Inputs into Embeddings
Unlike the dual-tower structure of CLIP-based methods, VLMs incorporate a vision encoder, a vision-language connector, and an LLM, enabling inputs to be flexibly specified as unimodal (image or text) or bimodal (image-text pairs) within a shared representation space. The visual input streamline is processed by the vision encoder followed by the vision-language connector, both of which are components of the VLM, to produce global context vision embeddings . Similarly, the textual streamline obtains textual embeddings through the text embedding layers of the LLM. To adapt the model to diverse downstream tasks, the query input further includes task definitions (e.g., “represent the given image with the following question”) between the image and text, where the task-specific instructions are drawn from MMEB (Jiang et al., 2025b) as well as our proposed SCaR.
4.2 Enabling Visual Prompts to VLM Embedders
While existing VLM-based embedding models establish a general mechanism to accommodate diverse modality combinations, they often struggle to incorporate visual interaction prompts and may fail to capture fine-grained entity-level information. A common workaround is to crop bounding boxes, but cropping yields coarse rectangular regions and ineffective results (see Appendix E.1) that may include background, ignore inter-entity relations (e.g., “next to”), or cut across multiple entities. To address these limitations, we propose the visual prompt streamline, which supports bounding boxes, clicks (points), and masks as inputs. The visual prompt is then processed with the given image by the segmentation model to produce a segmentation map, which is subsequently passed through our proposed segmentation-language connector to generate segmentation embeddings.
Segmentation Model. We adopt the pretrained SAM-2 (Ravi et al., 2025) as the segmentation model within VIRTUE. A key advantage of using SAM-2 is that treating segmentation as an entity-level feature instead of relying on cropping provides a structured prior that aligns with human perception of discrete objects, producing features that more faithfully capture the semantics of the referenced entity. Specifically, SAM-2 includes a prompt encoder, an image encoder, and a mask decoder333Visual prompts can be positive/negative for SAM-2, but we focus on positive prompts in this paper. We also omit the memory bank and memory attention designed for videos, as we focus on text and images.. The image encoder processes the visual input to produce unconditioned feature embeddings, whereas the prompt encoder accepts visual prompts to define the extent of the object in an image. To cover conventional non-visual-interactive scenarios (e.g., MMEB), we set the visual prompt to uniformly sampled points when it is not explicitly provided to leverage SAM-2’s inherent capability for automatic segmentation. This serves as a surrogate for user interactions and provides fine-grained entity-level cues in addition to the global context captured by the visual and textual streamlines. The mask decoder takes the outputs of the prompt encoder and the image encoder to produce a segmentation feature map . The mask prediction, IoU, occlusion MLP heads, and upsampling process in SAM-2’s mask decoder are discarded to utilize the segmentation feature map conditioned on visual prompts, since the segmentation feature map already encodes entity-level information from both the prompt and vision encoders; there is no need to merge multiple segmentation masks into a joint embedding space.
Segmentation-Language Connector. While we utilize the segmentation feature map to avoid the extra overhead of converting reconstructed masks into embeddings, directly flattening results in a sequence length of 4096, which cannot be feasibly processed when aligning to the LLM dimension due to GPU memory limitations. Therefore, we employ a 2D convolution layer to compress the feature map from 4096 tokens to . The resulting representations are then projected via two MLP layers: first into , and then into the LLM’s hidden dimension for joint learning as:
4.3 Training Schema
The segmentation , vision , and text embeddings are concatenated in the order of segmentation-vision-text and subsequently fed into the LLM to generate the query embedding or target embedding from the final hidden state of the last token for contrastive learning. In this manner, the query embedding is encouraged to move closer to semantically similar targets while being pushed away from dissimilar ones, incorporating not only global matching signals but also entity-level information provided by the segmentation embeddings.
Contrastive Learning. Since VLMs are not originally tailored for representation learning, we adopt contrastive learning with the InfoNCE loss (van den Oord et al., 2018) on query embeddings and target embeddings , each of which contains any combinations of , , and . Formally, InfoNCE is applied over in-batch negatives:
where denotes the set of hard negatives and is a temperature parameter. Following (Jiang et al., 2025b), GradCache (Gao et al., 2021) is employed to enable larger batch sizes, thereby improving the generalizability of the learned embeddings by leveraging a greater number of in-batch negatives.
5 Experiments
5.1 Experimental Setup
Implementation Details. We use Qwen2-VL-2B and Qwen2-VL-7B (Wang et al., 2024) as backbone VLMs for VIRTUE-2B and VIRTUE-7B, both with sam2.1_hiera_base_plus444https://github.com/facebookresearch/sam2 as the segmentation model. To adapt the pretrained VLMs for embedding tasks, we apply LoRA (Hu et al., 2022) to the LLM within VIRTUE following (Jiang et al., 2025b), while training the segmentation-language connector from scratch. The segmentation embeddings are prepended to the inputs only when images are provided as inputs. The segmentation model, vision encoder, and vision-language connector are kept frozen to preserve the pretrained knowledge. VIRTUE is trained with 20 in-distribution MMEB-train datasets with a batch size of 1024 and LoRA rank of 8. Detailed settings are provided in Appendix C, with ablation and parameter studies in Appendices E.1 and E.2.
Benchmarks. We evaluate VIRTUE across 20 in-distribution test sets and 16 out-of-distribution test sets from MMEB (Jiang et al., 2025b) to assess its universal instruction-following embedding capabilities across diverse classification, VQA, retrieval, and visual grounding scenarios. To examine the visual-interactive capability, VIRTUE is evaluated on our proposed SCaR benchmark comprising five datasets with both out-of-domain and in-domain performance. Consistent with the MMEB benchmark, we report precision@1 across all experiments.
Baselines. We strive to provide a large number of comparisons against recent CLIP-based and VLM-based families. As CLIP-based methods, we compare against CLIP (Radford et al., 2021), BLIP2 (Li et al., 2023), SigLIP (Zhai et al., 2023), OpenCLIP (Cherti et al., 2023), UniIR (Wei et al., 2024), and Magiclens (Zhang et al., 2024). As VLM-based methods, we compare against E5-V (Jiang et al., 2024), GME (Zhang et al., 2025), MMRet (Zhou et al., 2025), LamRA (Liu et al., 2025b), VLM2Vec (Jiang et al., 2025b), and UniME (Gu et al., 2025). For fair comparisons, results on MMEB are reported from their corresponding papers, and all models on SCaR, except for +SCaR-train, are used off-the-shelf, with frozen weights. GME, LamRA, and VLM2Vec adopt 2B and 7B of Qwen2-VL (Wang et al., 2024), E5-V uses LLaVA-NeXT-8B (Liu et al., 2024b), and MMRet and UniME use LLaVA-1.6-7B (Liu et al., 2024a).
| Per Meta-Task Score | Average Score | ||||||
| Classification | VQA | Retrieval | Grounding | IND | OOD | Overall | |
| Model #Datasets | 10 | 10 | 12 | 4 | 20 | 16 | 36 |
| w/o Finetuning on MMEB-Train | |||||||
| CLIPL/14(Radford et al., 2021) | 42.8 | 9.1 | 53.0 | 51.8 | 37.1 | 38.7 | 37.8 |
| BLIP2opt-2.7b (Li et al., 2023) | 27.0 | 4.2 | 33.9 | 47.0 | 25.3 | 25.1 | 25.2 |
| SigLIPso400m-14-384 (Zhai et al., 2023) | 40.3 | 8.4 | 31.6 | 59.5 | 32.3 | 38.0 | 34.8 |
| OpenCLIPL/14 (Cherti et al., 2023) | 47.8 | 10.9 | 52.3 | 53.3 | 39.3 | 40.2 | 39.7 |
| UniIRCLIP_CF (Wei et al., 2024) | 44.3 | 16.2 | 61.8 | 65.3 | 47.1 | 41.7 | 44.7 |
| MagiclensCLIP-L (Zhang et al., 2024) | 38.8 | 8.3 | 35.4 | 26.0 | 31.0 | 23.7 | 27.8 |
| GME-2B (Zhang et al., 2025) | 54.4 | 29.9 | 66.9 | 55.5 | 49.2 | 55.2 | 51.9 |
| E5-V-8B (Jiang et al., 2024) | 21.8 | 4.9 | 11.5 | 19.0 | 14.9 | 11.5 | 13.3 |
| GME-7B (Zhang et al., 2025) | 57.7 | 34.7 | 71.2 | 59.3 | 53.6 | 58.8 | 56.0 |
| LamRA-7B (Liu et al., 2025b) | 59.2 | 26.5 | 70.0 | 62.7 | 53.0 | 55.4 | 54.1 |
| w/ Finetuning on MMEB-Train | |||||||
| CLIPL/14 (Radford et al., 2021) | 55.2 | 19.7 | 53.2 | 62.2 | 47.6 | 42.8 | 45.4 |
| OpenCLIPL/14 (Cherti et al., 2023) | 56.0 | 21.9 | 55.4 | 64.1 | 50.5 | 43.1 | 47.2 |
| VLM2Vec-2B (Jiang et al., 2025b) | 58.7 | 49.3 | 65.0 | 72.9 | 64.9 | 53.3 | 59.7 |
| MMRet-7B (Zhou et al., 2025) | 56.0 | 57.4 | 69.9 | 83.6 | 68.0 | 59.1 | 64.1 |
| VLM2Vec-7B (Jiang et al., 2025b) | 62.7 | 56.9 | 69.4 | 82.2 | 71.4 | 58.1 | 65.5 |
| UniME-7B (Gu et al., 2025) | 60.6 | 52.9 | 67.9 | 85.1 | 68.4 | 57.9 | 66.6 |
| VIRTUE (Ours) w/ Finetuning on MMEB-Train | |||||||
| VIRTUE-2B | 64.1 | 55.7 | 68.4 | 78.7 | 69.7 | 58.8 | 64.8 |
| VIRTUE-7B | 65.6 | 60.4 | 71.8 | 87.3 | 74.4 | 61.4 | 68.6 |
| Improvements (2B) | +5.4 | +6.4 | +1.5 | +5.8 | +4.8 | +3.6 | +5.1 |
| Improvements (7B) | +2.9 | +3.0 | +0.6 | +2.2 | +3.0 | +2.6 | +2.0 |
5.2 Comparison on MMEB for Universal Embedding Tasks
Tab. 2 summarizes the overall performance of all methods, both with and without finetuning on MMEB-train. Our VIRTUE family consistently outperforms all baselines across the four core meta-task scenarios, as well as in both in-distribution (IND) and out-of-distribution (OOD) settings. Quantitatively, VIRTUE-2B achieves an average improvement of 5.1 points over CLIP-based and other 2B models (from 59.7 to 64.8), while VIRTUE-7B surpasses existing 7B models with a 2.0-point gain (from 66.6 to 68.6). While VLM-based models are superior to CLIP-based ones, VIRTUE significantly outperforms GME and LamRA with the same Qwen2-VL backbone, as well as MMRet and E5-V with different backbones across all meta-tasks, demonstrating the effectiveness and universality of VIRTUE. In addition, the comparisons between VIRTUE and VLM2Vec highlight the importance of incorporating entity-level information, since both models adopt the same training schema and data. The use of uniformly sampled points for non-visual-interactive tasks signifies that segmentation embeddings contribute positively to universal embedding performance by enriching global context with fine-grained object-level details. Compared to UniME-7B, VIRTUE-7B surpasses all meta-tasks scenarios, underscoring the generalized embedder ability of VIRTUE, which jointly integrates global context and segmentation-derived context with in-batch negatives.
5.3 Comparison on SCaR for Visual-Interactive Tasks
| Model Datasets | RefCOCO+ | RefCOCOg | VisualGenome | COCO-Stuff | ADE20K | Overall |
| w/o Finetuning on MMEB-Train | ||||||
| CLIP | 18.1 | 22.4 | 23.0 | 15.5 | 18.9 | 19.6 |
| ReCLIP | 17.6 | 23.0 | 21.3 | 12.6 | 15.9 | 18.1 |
| + Red Circle | 17.6 | 21.6 | 22.3 | 15.1 | 18.4 | 19.0 |
| GME-2B | 5.4 | 8.3 | 7.5 | 5.8 | 6.2 | 6.6 |
| +Cropping | 6.1 | 8.3 | 5.8 | 6.2 | 7.5 | 6.8 |
| GME-7B | 6.1 | 10.1 | 7.4 | 7.0 | 5.6 | 7.2 |
| +Cropping | 6.0 | 9.8 | 4.4 | 7.0 | 5.3 | 6.5 |
| LamRA-7B | 8.0 | 7.9 | 5.3 | 3.1 | 8.9 | 6.6 |
| +Cropping | 8.2 | 7.9 | 3.1 | 8.9 | 5.3 | 6.7 |
| w/ Finetuning on MMEB-Train | ||||||
| VLM2Vec-2B | 24.5 | 29.5 | 22.3 | 19.4 | 24.6 | 24.1 |
| +Cropping | 22.3 | 25.8 | 17.1 | 19.5 | 22.5 | 21.4 |
| +SCaR-train | 59.8 | 55.5 | 34.1 | 40.8 | 43.2 | 46.7 |
| VIRTUE-2B (Ours) | 28.8 | 42.4 | 24.4 | 29.9 | 27.5 | 30.4 |
| +Cropping | 24.4 | 36.3 | 14.2 | 25.8 | 23.4 | 24.8 |
| +SCaR-train | 64.2 | 65.3 | 41.4 | 54.2 | 56.0 | 56.2 |
| (2B) | +4.3 | +12.9 | +2.1 | +10.5 | +2.9 | +6.3 |
| (2B, +SCaR-train) | +4.4 | +9.8 | +7.3 | +13.4 | +12.8 | +9.5 |
| VLM2Vec-7B | 23.2 | 29.1 | 14.7 | 25.0 | 22.4 | 22.9 |
| +Cropping | 17.6 | 25.7 | 15.2 | 13.8 | 21.3 | 18.7 |
| +SCaR-train | 40.1 | 39.5 | 36.3 | 31.4 | 25.0 | 34.5 |
| MMRet-7B | 27.1 | 21.8 | 15.2 | 26.0 | 22.6 | 22.5 |
| +Cropping | 26.0 | 20.2 | 14.9 | 15.0 | 19.1 | 19.0 |
| +SCaR-train | 45.8 | 43.6 | 39.7 | 27.5 | 31.2 | 37.6 |
| UniME-7B | 31.4 | 32.8 | 19.0 | 25.3 | 23.0 | 26.3 |
| +Cropping | 25.8 | 24.0 | 15.2 | 24.6 | 22.3 | 22.4 |
| +SCaR-train | 57.8 | 59.3 | 44.5 | 41.9 | 43.6 | 49.4 |
| VIRTUE-7B (Ours) | 33.0 | 35.3 | 19.6 | 27.1 | 23.8 | 27.8 |
| +Cropping | 27.1 | 31.5 | 17.4 | 21.8 | 16.4 | 22.8 |
| +SCaR-train | 63.2 | 66.1 | 48.2 | 52.2 | 54.9 | 56.9 |
| (7B) | +1.6 | +2.5 | +0.6 | +1.8 | +0.8 | +1.5 |
| (7B, +SCaR-train) | +5.3 | +6.8 | +3.7 | +10.3 | +11.3 | +7.5 |
To examine the visual-interactive capabilities, VIRTUE is compared with CLIP as well as the crop-based (ReCLIP (Subramanian et al., 2022)) and visual-hinting-based (explicitly add red circles to images (Shtedritski et al., 2023), denoted as +Red Circle) variants. VLM-based embedding models are compared with and without finetuning on MMEB-train, due to their effectiveness and task-following abilities. Since none of the baselines naturally accept bounding boxes as visual inputs, they are textualized as Referring object bbox: {bbox}, including after queries, while VIRTUE takes not only the textualized input but also visual prompts. In addition, we further conduct experiments by naively cropping the specified object using the corresponding bounding box as the visual input for all models (+Cropping); in this case, no textualized bounding box is required. To push the limits of visual-interactive capabilities and for fair comparisons, we further finetune VIRTUE, VLM2Vec, MMRet, and UniME on the MMEB-train checkpoints for 1k steps using a batch size of 1024 and a learning rate of 2e-6, following the aforementioned formats, which are denoted as +SCaR-train.
Tab. 3 reports the results under both out-of-domain and in-domain (i.e., +SCaR-train) scenarios. Without finetuning on MMEB-train, CLIP outperforms VLM-based models. In contrast, simply adding visual hints, which was observed effectively in (Shtedritski et al., 2023), performs slightly worse than CLIP, likely due to the more challenging nature of SCaR, where captions emphasize reasoning and compositionality. VLM-based models without finetuning from MMEB-train (i.e., GME, LamRA) are more prone to underperform compared to models with finetuning from MMEB-train (i.e., VLM2Vec, MMRet, UniME), which can be attributed to the visual grounding datasets within MMEB-train that crop the ground-truth region as the target. Meanwhile, our proposed VIRTUE improves by 6.3 and 1.5 points on average for the 2B and 7B models, respectively, which implies that conditioning embeddings on visual prompts facilitates the learning of fine-grained representations for regions of interest. Although directly cropping the specified object enables a fine-grained understanding, the detrimental effects of +Cropping across all models and ReCLIP verify that it sacrifices global scene information for compositional reasoning. The results of +SCaR-train show that further finetuning with our collected SCaR training set is able to boost visual-interactive reasoning even for models that do not originally support visual prompts. Nonetheless, VIRTUE, equipped with the segmentation streamline, leads to a larger performance gain of 9.5 points for the 2B model and 7.5 points for the 7B model, which demonstrates the effectiveness of SCaR-train and incorporating visual prompts for not only fine-grained information but also user-enabled visual interactions. We include additional reevaluations on MMEB and qualitative results on SCaR with SCaR-train in Appendices E.3 and E.5, respectively. More application paradigms are presented in Appendix E.4.
6 Conclusion
In this paper, we proposed VIRTUE, a novel visual-interactive text-image universal embedder for both instruction-following and visual-interactive embedding scenarios. Distinct from existing models that support only text as the interaction modality, VIRTUE equipped with pre-trained SAM2 as the segmentation model is able to process visual prompts as inputs, enabling the model to jointly capture global context as well as entity-level information. We constructed SCaR, a large-scale segmentation-and-scene image-to-text retrieval benchmark, to assess VIRTUE’s visual-interaction reasoning abilities, which consists of 1M samples with challenging negative candidates generated by replacing elements of the ground-truth captions. Experiments on both MMEB and SCaR demonstrate that VIRTUE is consistently superior to state-of-the-art approaches by 3.1%-8.5% and 15.2%-20.3%, respectively. We believe VIRTUE serves as a generic framework for conventional instruction-following and visual-interactive embedding tasks, and that SCaR opens up new human-AI interaction opportunities for embedding models. We present further discussion of limitations and broader impacts in Appendices A and B.
Ethics Statements
The SCaR dataset is constructed from publicly available datasets (RefCOCO+, RefCOCOg, VisualGenome, COCO-Stuff, and ADE20K), which mitigates potential privacy concerns and harmful content. The dataset is primarily generated using GPT-4V with subsequent human inspection to remove ethically problematic content. While GPT-4V may inherit biases from its training data, our prompt design explicitly encourages diverse and creative captions, and a comprehensive filtering pipeline is applied to examine the generated captions. This design reduces human annotation bias and avoids repetitive lexical overlap, as illustrated in Fig. 6. Moreover, although VIRTUE builds upon a pre-trained VLM, it is further fine-tuned into an embedding model, which significantly minimizes the risk of harmful or copyrighted content generation while preserving the semantic richness needed for downstream tasks.
LLM Usage. In addition to GPT-4V used for building the SCaR benchmark, we use LLM solely for polishing the manuscript.
Reproducibility Statements
The codebase and partial SCaR benchmark samples (due to file size) are provided in the supplementary materials for reproducibility, and the experimental details as well as model configurations are summarized in Sec. 5.1 and Appendix C. VIRTUE-2B and VIRTUE-7B will be released in the camera-ready paper to advance the community with the realm of interactive representation learning. Evaluations on the MMEB benchmark follows their official use, while SCaR is detailed in Sec. 3 and Appendix D.
Appendix A Limitations and Future Directions
While our approach represents a significant step toward reinforcing visual-interactive capabilities in embedding models, the main limitation lies in training VIRTUE only with MMEB and SCaR, primarily due to computational constraints. Prior work (e.g., (Zhou et al., 2025)) has shown that incorporating more diverse datasets can further improve universality, suggesting that expanding our training sources would yield additional performance gains. Additionally, our evaluation emphasizes interactive image-to-text retrieval (via SCaR) as a primary validation task. While this provides strong evidence for the effectiveness of our approach, we only cover interactive image-to-image retrieval as case studies in Sec. E.4.1, largely due to copyright and ethical concerns associated with constructing suitable benchmarks. Nevertheless, this remains an important and complementary dimension for assessing visual-interactive embedding models. These limitations reflect our overarching objective: to pave the way for embedding models that are not only text-interactive but also visually interactive, while maintaining a strong universal embedding performance. In future work, we aim to extend our framework with more advanced training strategies and diverse as well as multifaceted datasets, including safe and ethically curated setups for interactive image-to-image retrieval, to fully unlock the potential of visual-interactive embedding models.
Appendix B Broader Impacts
As VIRTUE introduces a new paradigm for visual-interactive embedding models, it has the potential to serve as a multimodal encoder that naturally supports visual prompts for VLMs, in contrast to post-hoc finetuning approaches (e.g., (You et al., 2024; Lian et al., 2025)). It may also benefit the development of multimodal foundation models built on top of CLIP (e.g., video-text-to-audio generation models (Cheng et al., 2025), text-to-image generation models (Rombach et al., 2022)), given VIRTUE’s superior performance over existing CLIP-based methods. Moreover, embedding-based evaluation methods such as CLIPScore (Hessel et al., 2021) are widely used as automatic and reference-free metrics for holistic image-text alignment. VIRTUE could advance this line of evaluation by serving as a stronger embedding-based standard: while CLIPScore represents a subset functionality, VIRTUE further enables region-specific interactions in addition to capturing global context to compute text-image similarities.
Appendix C Implementation Details
| Configurations | VIRTUE-2B | VIRTUE-7B |
| MMEB-Train | ||
| VLM | Qwen2-VL-2B | Qwen2-VL-7B |
| Segmentation model | SAM2.1-hiera-base-plus | |
| Image resolution | 13441344 | |
| Segmentation-language connector | Random initialized | |
| LoRA rank | 8 | |
| LoRA dropout | 0.1 | |
| LoRA alpha | 64 | |
| Temperature | 0.02 | |
| Batch size | 1024 | |
| Training steps | 5000 | |
| Warmup steps | 200 | |
| Learning rate | 2e-5 | |
| 256 | ||
| 1536 | 3584 | |
| Sampled points | 9 | |
| # Segmentation embeddings | 256 | |
| GPU | 8H100 80G | |
| Precision | bf16 | |
| Optimizer | AdamW () | |
| Training time | 74 hours | 189 hours |
| SCaR-Train | ||
| VLM | Checkpoint of VIRTUE-2B | Checkpoint of VIRUTE-7B |
| Segmentation model | SAM2.1-hiera-base-plus | |
| Image resolution | 13441344 | |
| Segmentation-language connector | Checkpoint of VIRTUE-2B | Checkpoint of VIRUTE-7B |
| LoRA rank | 8 | |
| LoRA dropout | 0.1 | |
| LoRA alpha | 64 | |
| Temperature | 0.02 | |
| Batch size | 1024 | |
| Training steps | 1000 | |
| Warmup steps | 100 | |
| Learning rate | 2e-6 | |
| 256 | ||
| 1536 | 3584 | |
| Sampled points | 9 | |
| # Segmentation embeddings | 256 | |
| GPU | 8H100 80G | |
| Precision | bf16 | |
| Optimizer | AdamW () | |
| Training time | 12 hours | 30 hours |
VIRTUE-2B and VIRTUE-7B are trained with the of 0.02 and learning rate of 2e-5, and input resolution of . The lengths of vision and text embeddings default to VLMs. Since segmentation embeddings are prepended to the sequence, their attention masks and positional encodings are set sequentially, following the original VLM design. is 256 from SAM2.1’s default configuration, and is 1536 for 2B and 3584 for 7B from Qwen2-VL’s default settings. For evaluating MMEB, VIRTUE is trained with MMEB-train with 5k steps, and is further trained with 1k steps with SCaR-train for (+SCaR-train in Tab. 3). The VIRTUE 2B and 7B training was conducted on 8H100 80GB, where MMEB-train took around 74 and 189 hours, respectively, and SCaR-train took around 12 and 30 hours, respectively. The kernel size and stride of , the number of sampled points , and the number of segmentation embeddings are empirically set to 4, 9, and 256, respectively. Detailed configurations are summarized in Tab. 4.
Appendix D SCaR Details
D.1 Dataset Details
We summarize the five datasets used for building SCaR below:
RefCOCO+ (Yu et al., 2016): A large-scale dataset for referring expression comprehension and segmentation, containing around 45k expressions over 19k images from MS-COCO. Crucially, it prohibits location-based words (e.g., “left”, “right”), forcing models to rely on appearance and contextual reasoning–a more robust form of vision-language grounding that moves beyond simple spatial relationships. Its unique constraints make it a strong test of a model’s ability to truly understand an object’s visual properties relative to its surroundings.
RefCOCOg (Mao et al., 2016): A referring expression dataset with around 50k expressions for around 30k images, collected with longer and more descriptive annotations. This rich linguistic detail makes it a standard benchmark for evaluating fine-grained vision-language grounding, as it requires models to reason over nuanced and complex natural language descriptions rather than just a few keywords.
VisualGenome (Krishna et al., 2017): A massive vision-language dataset that provides dense annotations for over 100k images, including objects, attributes, and region-level descriptions. Its most notable feature is the inclusion of scene graph annotations, which explicitly model relationships between objects. This structured data is invaluable for training and evaluating models on compositional reasoning, enabling a deeper understanding of complex scenes beyond simple object detection.
COCO-Stuff (Caesar et al., 2018): An extension of the MS-COCO dataset with over 160k images annotated for 91 “stuff” classes (e.g., sky, grass) in addition to the existing “thing” categories. This comprehensive annotation scheme makes it a primary benchmark for semantic segmentation and tasks that require holistic scene understanding, as it allows models to learn the fine-grained contextual relationships between foreground objects and their background environments.
ADE20K (Zhou et al., 2017): A challenging scene parsing dataset with 20k images and 150 fine-grained semantic categories. The dataset’s diversity, spanning a wide range of indoor and outdoor scenes, coupled with its high-quality pixel-level labels, makes it an essential standard for evaluating the performance of semantic segmentation and general scene understanding models.
D.2 Prompt Template for Building SCaR
Fig. 4 illustrates the detailed prompt template used for building the SCaR dataset (Sec. 3.2), where each sample varies in image size, caption, category, and bounding box (bbox) as provided in the original datasets. The ground-truth instruction guides GPT-4V to determine whether the three elements <object> <relation> <scene> are satisfied; if any element is missing, GPT-4V is required to complete it. Subsequently, the negatives instruction directs GPT-4V to replace each element from the ground-truth caption. Notably, we explicitly include creativity and diversity instructions to encourage GPT-4V to generate more diverse and imaginative negatives, as visualized in the word clouds in Fig. 6. To ensure GPT-4V follows these instructions, we provide an example output at the end of the prompt.
D.3 Prompt Template for LLM-based Filtering
Fig. 5 shows the prompt for LLM-based filtering. Generally, we adopt GPT-4V to guide the data verification from multifaceted perspectives. If all conditions are passed, the output JSON object contains the status field with “passed” as well as the extracted objects, relations, and scenes for both ground-truth and negative captions, which are then passed on to WordNet to inspect again in terms of semantic differences.
D.4 Additional SCaR Statistics
We further analyze the composition of SCaR to better understand its benchmark characteristics. Fig. 6 summarizes detailed statistics, including word clouds, annotation counts, and image distributions for both the training and evaluation splits. Attributed to the creative and diverse instruction collection process described in Sec. 3.2, we observe that the word frequency distributions of ground-truth and negative captions are closely aligned across splits, suggesting that the benchmark is not dominated by a small set of frequent words. Moreover, both image- and annotation-level distributions reveal that COCO-Stuff and VisualGenome contribute the largest proportions of samples, reflecting their broad coverage of scenes and object relationships. Fig. 7 illustrates the sentence length distributions of ground-truth and negative captions across the five datasets in SCaR. Overall, the two distributions are closely aligned, indicating that the synthesized negative captions generated by GPT-4V preserve similar linguistic complexity to the ground-truth annotations rather than introducing trivial artifacts. We also observe slightly heavier tails in the negative caption distributions, suggesting greater variability in sentence length. This property enhances the linguistic diversity within SCaR while maintaining a comparable level of difficulty, thereby preventing models from exploiting superficial cues such as caption length.
D.5 SCaR Examples
| Dataset | Query Image with bbox | Candidates | ||||||||||
| RefCOCO+ (Yu et al., 2016) | ![]() |
| RefCOCOg (Mao et al., 2016) | ![]() |
| VisualGenome (Krishna et al., 2017) | ![]() |
| COCO-Stuff (Caesar et al., 2018) | ![]() |
| ADE20K (Zhou et al., 2017) | ![]() |
Tab. 5 presents SCaR examples for each dataset, where it can be seen that the five datasets cover diverse domains. Additionally, the examples indicate that naively cropping with bounding boxes would sacrifice global scene context (e.g., “at the beach” in RefCOCOg, “by the waterfront” in VisualGenome, and “in a glass-roofed lounge” in ADE20K).
Appendix E Extensive Experiments
E.1 Ablation Study
| 1) Choice | MMEB | SCaR |
| Phi-3.5-V | 61.7 | 26.5 |
| Qwen2-VL-2B | 64.8 | 30.4 |
| 2) Instruction | MMEB | SCaR |
| ✗ | 59.2 | 24.3 |
| ✓ | 64.8 | 30.4 |
| 3) Resolution | MMEB | SCaR |
| 672x672 | 60.1 | 27.6 |
| 1344x1344 | 64.8 | 30.4 |
| 4) Alternative | MMEB | SCaR |
| Prompter | 61.0 | 22.7 |
| Cropped | 63.3 | 25.9 |
| Segmentation | 64.8 | 30.4 |
| 5) Choice | MMEB | SCaR |
| SAM2.1-S | 60.1 | 21.6 |
| SAM2.1-B+ | 64.8 | 30.4 |
| SAM2.1-L | 63.8 | 27.9 |
| 6) # MLP | MMEB | SCaR |
| 1 | 60.4 | 18.6 |
| 2 | 64.8 | 30.4 |
| 3 | 63.5 | 29.7 |
| 7) | MMEB | SCaR |
| 64 | 61.1 | 29.9 |
| 256 | 64.8 | 30.4 |
| 1024 | 60.2 | 35.1 |
We conduct seven ablation variants to investigate the relative contributions of different components: 1) VLM backbones, 2) with and without an instruction in the query, 3) input image resolution, 4) segmentation streamline alternatives, 5) different sizes of SAM-2, and 6) different numbers of MLPs in the segmentation-language connector (Eq. 1), where one MLP projects from to , and three MLPs project from to 768, then to 1024, and finally to ; and 7) segmentation embedding lengths . For 4) segmentation streamline alternatives, we replace our segmentation model with Prompter from CLOC555CLOC cannot be directly evaluated or reproduced, as its model, code, and parts of training data are not publicly available. We attempted to contact the authors but did not receive a response. (Chen et al., 2025), which encodes bounding boxes via a randomly initialized Transformer, or directly crops images for the vision encoder (denoted as Cropped). Both variants concatenate embeddings from nine random crops, similar to the setting of nine sampled points in SAM2.1 for the segmentation-language connector for MMEB.
As shown in Tab. 6, although Phi-3.5-V has more parameters than Qwen2-VL-2B, the latter yields better results on MMEB and SCaR. Removing text instructions and using lower resolutions lead to a significantly inferior performance. The degraded results from Prompter and cropping suggest that pre-trained SAM2.1 provides not only more precise segmentations than simple crops but also a stronger understanding from its pre-trained knowledge. Although SAM2.1-L performs on par with SAM2.1-B+ in terms of MMEB, its performance degrades on SCaR. In contrast, using SAM2.1-S deleteriously impacts both benchmarks. Additionally, increasing MLP layers and segmentation embedding length does not consistently yield gains across MMEB and SCaR. While increasing segmentation embeddings to 1024 delivers better results on SCaR, we stick to selecting 256 as our configuration to balance computation and effectiveness.
E.2 Hyperparameter Study
We further analyze the impacts of 1) LoRA rank (4, 8, 16, 32), 2) batch size (256, 512, 1024, 2048), and 3) sampling points used in non-visual-interactive tasks (MMEB). Fig. 8 shows that varying LoRA ranks produces consistent outcomes, with rank 8 showing a slight edge. Furthermore, larger batch sizes improve generalizability for in-batch negatives, with a batch size of 1024 obtaining the best results. We also find that increasing the number of sampling points does not enhance model effectiveness for non-visual-interactive tasks, presumably because key entities within an image for general tasks can be captured with only a few points.
E.3 Reevaluating MMEB with Further Finetuning on SCaR-train
| Per Meta-Task Score | Average Score | ||||||
| Classification | VQA | Retrieval | Grounding | IND | OOD | Overall | |
| # Datasets | 10 | 10 | 12 | 4 | 20 | 16 | 36 |
| VLM2Vec-2B | 58.7 | 49.3 | 65.0 | 72.9 | 64.9 | 53.3 | 59.7 |
| +SCaR-train | 35.8 | 7.5 | 14.7 | 25.8 | 19.4 | 19.2 | 19.8 |
| VLM2Vec-7B | 62.7 | 56.9 | 69.4 | 82.2 | 71.4 | 58.1 | 65.5 |
| +SCaR-train | 47.5 | 31.8 | 48.6 | 46.3 | 44.4 | 42.0 | 43.4 |
| MMRet-7B | 56.0 | 57.4 | 69.9 | 83.6 | 68.0 | 59.1 | 64.1 |
| +SCaR-train | 45.2 | 38.1 | 51.7 | 77.9 | 51.8 | 46.5 | 49.4 |
| UniME-7B | 60.6 | 52.9 | 67.9 | 85.1 | 68.4 | 57.9 | 66.6 |
| +SCaR-train | 55.0 | 46.4 | 64.3 | 84.5 | 62.7 | 54.3 | 59.0 |
| VIRTUE-2B (Ours) | 64.1 | 55.7 | 68.4 | 78.7 | 69.7 | 58.8 | 64.8 |
| +SCaR-train | 64.7 | 55.1 | 68.3 | 79.2 | 69.8 | 58.7 | 64.9 |
| VIRTUE-7B (Ours) | 65.6 | 60.4 | 71.8 | 87.3 | 74.4 | 61.4 | 68.6 |
| +SCaR-train | 64.4 | 58.1 | 67.0 | 86.4 | 72.2 | 60.0 | 66.8 |
To evaluate the impact of incorporating SCaR-train on MMEB performance, we reevaluate VLM2Vec-2B, VLM2Vec-7B, MMRet-7B, UniME-7B, VIRTUE-2B, and VIRTUE-7B with SCaR-train (+SCaR-train) (Sec. 5.3) and compare them against the original checkpoints trained solely on MMEB-train. As shown in Tab. 7, all baselines exhibit a degraded performance, likely because the models overfit or shift focus toward the SCaR distribution for incremental learning (Li & Hoiem, 2016). Notably, VLM2Vec and MMRet suffer severe degradation, suggesting stronger susceptibility to catastrophic forgetting (French, 1999) when their training relies more heavily on additional data. In contrast, UniME, with its multi-stage contrastive learning design, demonstrates greater robustness against forgetting. Meanwhile, VIRTUE maintains a performance comparable to the original MMEB-trained checkpoints, with VIRTUE-2B even achieving a slight 0.1-point gain. We attribute this stability to the segmentation streamline, which enables the use of optional visual prompts and makes VIRTUE more adaptable as a universal embedder capable of handling both visual-interactive and non-visual-interactive tasks.
E.4 Case Studies for VIRTUE’s Visual-Interactive Capabilities
With visual interaction, VIRTUE enables new applications such as segment-level retrieval, where users select a region of interest to fetch semantically matching images, and entity-level hinting for on-the-fly correction, thereby extending the utility of embedding-based systems far beyond traditional global matching.
E.4.1 Visual-Interactive In-the-Wild I2I Retrieval
While evaluating VIRTUE on SCaR confirms its ability to interact with visual interactions, we further conduct in-the-wild visual-interactive image-to-image retrieval. Similar to the procedure we constructed SCaR, we prompt GPT-4o with “Generate some text captions that need to consist of “object, relation, scene” for searching images. Only contain 2 entities within an image is sufficient.”. Then, we use the generated prompt to the Google Search API666https://developers.google.com/custom-search/ to get six ground-truth images (three depicting cars parked on the road and three depicting dogs on the sidewalk with nearby grass), and subsequently instruct GPT-4o to generate 20 negative captions conditioned on the selected text caption. Afterwards, those negative captions are used to search the corresponding images as negative candidates.
Fig. 9 illustrates the visual-interactive (top three rows) and naive cropping (bottom two rows) image-to-image paradigm enabled by VIRTUE-7B. The first row depicts conventional retrieval applications, also supported by VIRTUE, where images are retrieved based on holistic query image information. What distinguishes VIRTUE is its ability to incorporate user-specified visual prompts, as demonstrated in the second and third rows within the figure. For instance, selecting the parking car directs VIRTUE to focus on the car while preserving the surrounding scene context (e.g., grass or trees along the road). Similarly, selecting the dog on a sidewalk guides the model to retrieve dogs on sidewalks with nearby grass or trees. Attributed to the segmentation streamline in VIRTUE, both scenarios illustrate novel possibilities for interactive querying between humans and embedding models. On the flip side, naive cropping (bottom two rows) discards scene context, leading the model to retrieve the selected entity in mismatched contexts from the given image (e.g., a car running on the street or a dog running on the grass in the retrieved results, even though neither is moving in the query image).
E.4.2 On-the-Fly Correction with Visual Hinting
To delve into the capability of on-the-fly correction, we randomly sample MMEB cases that VIRTUE-2B initially misclassifies. We then manually provide the correct regions as visual prompts and re-run inference. As illustrated in Fig. 10, VIRTUE successfully corrects predictions not only for VQA but also for retrieval tasks, relying solely on visual hints at inference time without additional finetuning. This demonstrates a new mode of applicability for VLM-based embedding models, where users can guide the model interactively while avoiding the computational cost of conventional finetuning. Although direct cropping of hinted regions may also yield corrections, visual prompting in VIRTUE eliminates the need for heuristic preprocessing and proves more robust in challenging cases. For example, in VQA tasks, cropping can fail when the region excludes crucial contextual information (e.g., the motorcyclist), whereas visual prompts enable the model to integrate both global and fine-grained cues.
E.5 Qualitative Results on SCaR
Fig. 11 presents qualitative comparisons between baselines and our VIRTUE-7B model trained with SCaR-train. We observe that the baselines frequently misinterpret relational cues (e.g., “player with head down” to “player picking up a bat”) and even retrieve entirely incorrect objects (e.g., “tree” to “cow and statue”). A likely reason is that their representations are dominated by LLM-derived features (Tong et al., 2024), which tend to overlook spatial information such as bounding boxes in the text query. In contrast, VIRTUE incorporates a segmentation-based streamline that augments regions of interest with fine-grained visual representations, effectively guiding retrieval toward semantically and spatially accurate matches.
E.6 Detailed Scores of MMEB
Tab. 8 presents detailed results of 6 VLM-based models that are trained with MMEB for comprehensive comparisons. The performances are sourced from the corresponding papers. Due to the space limits, the detailed scores of CLIP, BLIP2, SigLIP, OpenCLIP, UniIR, Magiclens, E5-V, GME, and LamRA can be referred from their official papers.
| VLM2Vec-2B | VIRTUE-2B | MMRet-7B | VLM2Vec-7B | UniME-7B | VIRTUE-7B | |
| Classification (10 tasks) | ||||||
| ImageNet-1K | 77.5 | 80.1 | 58.1 | 80.1 | 71.3 | 82.3 |
| N24News | 73.7 | 78.4 | 71.3 | 79.7 | 79.5 | 81.1 |
| HatefulMemes | 58.3 | 67.5 | 53.7 | 69.7 | 64.6 | 74.1 |
| VOC2007 | 74.3 | 83.1 | 85.0 | 80.7 | 90.4 | 85.4 |
| SUN397 | 73.8 | 74.3 | 70.0 | 77.4 | 75.9 | 78.3 |
| Place365 | 35.3 | 36.4 | 43.0 | 37.4 | 45.6 | 39.6 |
| ImageNet-A | 50.9 | 53.4 | 36.1 | 58.1 | 45.5 | 55.2 |
| ImageNet-R | 84.7 | 88.3 | 71.6 | 73.9 | 78.4 | 82.5 |
| ObjectNet | 37.1 | 48.9 | 55.8 | 40.1 | 36.4 | 44.3 |
| Country-211 | 21.5 | 30.5 | 14.7 | 29.8 | 18.7 | 32.8 |
| Average | 58.7 | 64.1 | 56.0 | 62.7 | 60.6 | 65.6 |
| VQA (10 tasks) | ||||||
| OK-VQA | 48.5 | 56.7 | 73.3 | 56.8 | 68.3 | 59.6 |
| A-OKVQA | 39.5 | 50.5 | 56.7 | 47.3 | 58.7 | 52.2 |
| DocVQA | 82.5 | 87.9 | 78.5 | 89.7 | 67.6 | 91.7 |
| InfographicsVQA | 47.7 | 53.4 | 39.3 | 60.0 | 37.0 | 63.7 |
| ChartQA | 42.3 | 48.2 | 41.7 | 56.9 | 33.4 | 61.0 |
| Visual7W | 51.2 | 52.3 | 49.5 | 52.7 | 51.7 | 55.0 |
| ScienceQA | 30.7 | 40.9 | 45.2 | 38.5 | 40.5 | 46.7 |
| VizWiz | 38.6 | 44.3 | 51.7 | 39.9 | 42.7 | 43.3 |
| GQA | 48.3 | 48.7 | 59.0 | 55.1 | 63.6 | 54.7 |
| TextVQA | 63.3 | 74.1 | 79.0 | 71.6 | 65.2 | 76.4 |
| Average | 49.3 | 55.7 | 57.4 | 56.9 | 52.9 | 60.4 |
| Retrieval (12 tasks) | ||||||
| VisDial | 74.3 | 78.5 | 83.0 | 81.9 | 79.7 | 83.3 |
| CIRR | 46.8 | 56.0 | 61.4 | 51.1 | 52.2 | 60.4 |
| VisualNews_t2i | 73.1 | 75.1 | 74.2 | 80.5 | 74.8 | 80.8 |
| VisualNews_i2t | 73.7 | 77.9 | 78.1 | 81.2 | 78.8 | 82.5 |
| MSCOCO_t2i | 73.4 | 73.9 | 78.6 | 77.2 | 74.9 | 78.0 |
| MSCOCO_i2t | 68.5 | 72.9 | 72.4 | 73.9 | 73.8 | 76.3 |
| NIGHTS | 66.3 | 66.9 | 68.3 | 67.6 | 66.2 | 70.6 |
| WebQA | 85.9 | 88.7 | 90.2 | 88.3 | 89.8 | 91.2 |
| FashionIQ | 14.0 | 15.0 | 54.9 | 17.1 | 16.5 | 18.1 |
| Wiki-SS-NQ | 54.2 | 60.8 | 24.9 | 62.3 | 66.6 | 66.3 |
| OVEN | 68.3 | 69.2 | 87.5 | 66.5 | 55.7 | 67.2 |
| EDIS | 81.2 | 85.6 | 65.6 | 85.7 | 86.2 | 86.6 |
| Average | 65.0 | 68.4 | 69.9 | 69.4 | 67.9 | 71.8 |
| Visual Grounding (4 tasks) | ||||||
| MSCOCO | 66.5 | 70.7 | 76.8 | 76.5 | 76.5 | 80.5 |
| RefCOCO | 80.9 | 86.0 | 89.8 | 89.3 | 89.3 | 94.2 |
| RefCOCO-matching | 75.7 | 83.1 | 90.6 | 90.6 | 90.6 | 92.0 |
| Visual7W-pointing | 68.3 | 75.0 | 77.0 | 84.1 | 84.1 | 82.6 |
| Average | 72.9 | 78.7 | 83.6 | 82.2 | 85.1 | 87.3 |
| Final Score (36 tasks) | ||||||
| IND | 64.9 | 69.7 | 68.0 | 71.4 | 68.4 | 74.4 |
| OOD | 53.3 | 58.8 | 59.1 | 58.1 | 57.9 | 61.4 |
| Average | 59.7 | 64.8 | 64.1 | 65.5 | 66.6 | 68.6 |
![[Uncaptioned image]](Figures/scar-examples/refcoco_plus.png)
![[Uncaptioned image]](Figures/scar-examples/refcocog.png)
![[Uncaptioned image]](Figures/scar-examples/visualgenome.png)
![[Uncaptioned image]](Figures/scar-examples/coco-stuff.png)
![[Uncaptioned image]](Figures/scar-examples/ADE20K.jpg)